# Rethinking Misalignment in Vision-Language Model Adaptation from a Causal Perspective

### v1

@@ 5.4 Ablative Experiments
The effectiveness of multiple templates and DSTC. In Table 4, all results except the baseline method are obtained by increasing the number of templates to four. In the second row, we train the learnable tokens within the templates using the conventional cross-entropy loss and average the predictions from different templates to generate the final classification results. Compared to the baseline, multiple templates lead to a 0.68% improvement in the performance of base classes and a 0.92% enhancement in the performance of new classes, resulting in an overall HM gain of 0.81%. While modeling the uncertainty of each template, as shown in the third line, the performance for new classes achieves an improvement of 0.72%, and HM increases by 0.25%. Without explicitly decoupling the semantics within different templates, increasing the number of templates effectively boosts the performance. We argue that the semantics acquired by distinct templates, each with varied initialization, is inherently independent (See analysis in Appendix D). Therefore, simply applying multiple templates guided by the front-door adjustment could enhance the performance.
The effectiveness of VSD. The fourth and fifth rows of Table 4 present the experimental results obtained by decoupling semantics solely within the image and text branches, respectively. When compared to the third row, it becomes evident that decoupling semantics at either branch alone does not yield a substantial improvement in the Harmonic Mean (HM). However, upon comparing the sixth row with the third row, it is observed that when both decoupling methods are employed concurrently, the model achieves a 0.69% increase in base class performance and a 0.60% improvement in novel class performance, leading to an overall HM enhancement of 0.64%. Despite the inherent decoupling of semantics among different templates, the utilization of VSD further promotes the learning of distinct semantics for each template, thereby enhancing the model’s overall performance.
@@ 6 Conclusion and Future Discussion
Appendix A Algorithm
  repeat
     for $i$-$th$ training iteration do
        Iteratively sample a minibatch $\mathcal{X}^{\prime}$ from $\mathcal{X}$.
        for $m$-$th$ template do
           ${\mathcal{X}^{\prime}}^{m}=\mathcal{A}^{m}(\mathcal{X}^{\prime})$.
           Generate the text features $w^{m}$ with the text encoder.
           Using Equation 6 to generate the classification evidence $e^{m}$ with $v^{m}$ and $w^{m}$.
           Using Equation 9 to calculate $\mathcal{L}_{t\text{-}ce}$.
           Using Equation 3 and Equation 5 to calculate $\mathcal{L}_{de}$ and $\mathcal{L}_{con}$.
           Using Equation 10 to calculate $\mathcal{L}_{CDC}$.
           $\theta^{m}=\theta^{m}-\ell\nabla_{\theta}\mathcal{L}_{CDC}$.
        end for
     end for
  until $\theta$ converge.
Algorithm 1 The training pipeline of CDC
  Sample a test data $x$ from $X$
  for $m$-$th$ template do
     Generate the image features $v^{m}$ of $x$.
     Generate the text features $w^{m}$ with the text encoder.
     Using Equation 6 to generate the classification evidence $e^{m}$ with $v^{m}$ and $w^{m}$.
  end for
  Using Equation 8 to generate the final classification results.
Algorithm 2 The testing pipeline of CDC
Appendix B Hand-Crafted Prompts
@@
Appendix C Evaluation Protocol
@@
where Base and New denote the accuracy on base classes and new classes, respectively.
@@
Appendix D Additional Experiments
D.1 Analysis of the Hyper-Parameters
@@
(a) UCF101
(b) Caltech101
Figure 5: The impact of $\gamma$ on performance.
D.2 Template Study
@@
D.3 Intrinsic Decoupling of Semantics from Different Templates

### v2

@@ 5.4 Ablative Experiments
The effectiveness of multiple templates and DSTC. In Table 4, all results except the baseline method are obtained by increasing the number of templates to four. In the second row, we train the learnable tokens within the templates using the conventional cross-entropy loss and average the predictions from different templates to generate the final classification results. Compared to the baseline, multiple templates lead to a 0.68% improvement in the performance of base classes and a 0.92% enhancement in the performance of new classes, resulting in an overall HM gain of 0.81%. While modeling the uncertainty of each template, as shown in the third line, the performance for new classes achieves an improvement of 0.72%, and HM increases by 0.25%. Without explicitly decoupling the semantics within different templates, increasing the number of templates effectively boosts the performance. We argue that the semantics acquired by distinct templates, each with varied initialization, is inherently independent (See analysis in Appendix C). Therefore, simply applying multiple templates guided by the front-door adjustment could enhance the performance.
The effectiveness of VSD. The fourth and fifth rows of Table 4 present the experimental results obtained by decoupling semantics solely within the image and text branches, respectively. When compared to the third row, it becomes evident that decoupling semantics at either branch alone does not yield a substantial improvement in the accuracy of new classes. However, upon comparing the sixth row with the third row, it is observed that when both decoupling methods are employed concurrently, the model achieves a 0.69% increase in base class performance and a 0.60% improvement in novel class performance, leading to an overall HM enhancement of 0.64%. Despite the inherent decoupling of semantics among different templates, the utilization of VSD further promotes the learning of distinct semantics for each template, thereby enhancing the model’s overall performance.
@@ 6 Conclusion and Future Discussion
Acknowledgments
The authors gratefully acknowledge the valuable feedback provided by the anonymous reviewers. This work is supported by the Fundamental Research Program, China, Grant No.JCKY2022130C020.
Appendix A Hand-Crafted Prompts
@@
Appendix B Evaluation Protocol
@@
where Base and New denote the accuracy of base classes and new classes, respectively.
Appendix C Additional Experiments
C.1 Analysis of the Hyper-Parameters
@@
(a) UCF101
(b) Caltech101
Figure 5: The impact of $\gamma$ on performance.
@@
C.2 Template Study
@@
C.3 Intrinsic Decoupling of Semantics from Different Templates